From the circuit shown determine:

a) VCE, VGS, VDS

b) AV=V0=Vi

Data: \beta =100, IDSS=6mA, VP= -6V

220V 2.2K 40KS 0.1uF E ovo 0.1uF VaR3JFET 10K 1.2K
Answer #1

12ov. G puen ckt 82oak S40 lok & aka De conditpons Tau Rth Rth yoklliol = OXIO 40+10 = not . 59 | Rth=8kr КЛ. vth: Von XRGTGS - 1-60) Vos = VD-us. Vpa. Upis IORD= 20 - 2-729A AX2- 24 Pub=!4-02Ü VSEVGS= -1.6 V VDS= vip-us= 14.02- (-1-6)u. = 14.02 +Smali signal model . Gate brain is mugs $RD X 3Rth . इ०९ SRG 2 Vo= - Smugs RD sin = 220 Vigs=09 = -9mRD LanRi9n = 2ID = ;-1-3